Imagine you have some workers and some hand-held computers that you can use to take inventory at a warehouse. There are diminishing returns to taking inventory. If one worker uses one computer, he can inventory 125 items per hour. Two workers sharing a computer can together inventory 175 items per hour. Three workers sharing a computer can together inventory 225 items per hour. And four or more workers sharing a computer can together inventory 235 items per hour. Computers cost $125 each and you must pay each worker $30 per hour.

a. If you assign one worker per computer, what is the cost of inventorying a single item with one worker?

Instructions: Round your answer to two decimal places.

$ _____ per item

b. If you assign two workers per computer, what is the cost of inventorying a single item?

Instructions: Round your answer to two decimal places.

$ ______ per item

c. If you assign three workers per computer, what is the cost of inventorying a single item?

Instructions: Round your answer to two decimal places.

$ _____ per item

d. If you assign four workers per computer, what is the cost of inventorying a single item?

Instructions: Round your answer to two decimal places.

$ ______ per item

e. How many workers per computer should you assign if you wish to minimize the cost of inventorying a single item?

________ worker(s) per computer
Answer #1

ANSWER:

(a)

When one worker per computer is assigned then, in such case, worker can inventory 125 items per hour.

Total cost = Cost of computer + Wage of worker

Total cost = $125 + $30

Total cost = $155

Calculate the cost of inventorying a single item -

Cost = Total cost/Total items = $155/125 = $1.24

The cost of inventorying a single item is $1.24.

(b)

When two workers per computer is assigned then, in such case, workers can inventory 175 items per hour.

Total cost = Cost of computer + Wage of workers

Total cost = $125 + $60

Total cost = $185

Calculate the cost of inventorying a single item -

Cost = Total cost/Total items = $185/175 = $1.06

The cost of inventorying a single item is $1.06.

(c)

When three workers per computer is assigned then, in such case, workers can inventory 225 items per hour.

Total cost = Cost of computer + Wage of worker

Total cost = $125 + $90

Total cost = $215

Calculate the cost of inventorying a single item -

Cost = Total cost/Total items = $215/225 = $0.95

The cost of inventorying a single item is $0.95.

(d)

If the cost of inventorying a single item has to be minimized then 3 workers per computer should be

* Please rate the answer if you liked it by clicking on thumbs-Up. Thank-you